# Adaptive Group Robust Ensemble Knowledge Distillation

### v1

@@ Appendix C Supplemental experiments
Figure 4: Results on student model capacity. We perform experiments on WaterBirds and CelebA using resnet50 teachers and varying the capacity of the student network (resnet18, resnet34, and resnet50). We plot the average and the worst-case accuracy of different ensemble distillation methods across three random seeds. The WGA tends to increase as the student model has more capacity for learning the core features. Most baselines match or outperform the teachers when the student model has the same capacity as the teachers (i.e., resnet50), and our method remains superior in terms of WGA. These results suggest that the reduced capacity of the student model is a source of the disparity observed.
Figure 5: Effect of the temperature parameter on the WGA. We study the effect of the temperature parameter (Eq. 3) in WaterBirds and CelebA datasets using five resnet50 teachers and a resnet18 student. We perform knowledge distillation with different temperature parameters $\tau\in\{1,2,4,...,10\}$ and report the average test WGA across three independent random seeds. The WGA can improve with increased temperature and steady or decrease after certain higher values of temperature ($\tau\geq 10$ and $\tau\geq 8$ for the Waterbirds and CelebA datasets, respectively). These results align with recent work by Mohammadshahi and Ioannou, (2024) studying the benefit of increased temperature for fairness in knowledge distillation.
Table 3: Sensitivity of AGRE-KD to the biased model architecture. In the main paper, we used a biased model with the same architecture as the teacher models (i.e., resnet50). We experiment with different network backbones for the biased models in AGRE-KD. The results below do not show significant differences across biased model choices, demonstrating the robustness of the proposed method to the choice of biased model. These results suggest that the gradient direction of any biased pretrained model can provide sufficient guidance for debiased distillation.
@@
Resnet18 $90.2_{\pm 1.10}$ $86.8_{\pm 2.88}$ $91.5_{\pm 0.09}$ $89.8_{\pm 0.16}$
@@
Figure 6: Effect of teacher size. We train the resnet18 student model with different ensemble sizes of resnet50 teachers (5, 10, 20, …, 50). Increasing ensemble size exerts a positive effect on both the average and the worst-group performance. However, the Random KD method tends to get worse as we increase the number of teachers.

### v2

@@ Appendix D Supplemental results
Figure 6: Results on student model capacity. We perform experiments on WaterBirds and CelebA using ResNet-50 teachers and varying the capacity of the student network (ResNet-18, ResNet34, and ResNet-50). We plot the average and the worst-case accuracy of different ensemble distillation methods across three random seeds. The WGA tends to increase as the student model has more capacity for learning the core features. Most baselines match or outperform the teachers when the student model has the same capacity as the teachers (i.e., ResNet-50), and our method remains superior in terms of WGA. These results suggest that the reduced capacity of the student model is a source of the disparity observed.
Figure 7: Effect of ensemble size. We train the ResNet-18 student model with different ensemble sizes of ResNet-50 teachers (5, 10, 20, …, 50). Increasing ensemble size exerts a positive effect on both the average and the worst-group performance. However, the Random KD method tends to get worse as we increase the number of teachers.
@@
ResNet-18 $90.2_{\pm 1.10}$ $86.8_{\pm 2.88}$ $91.5_{\pm 0.09}$ $89.8_{\pm 0.16}$
Table 5: Sensitivity of AGRE-KD to the biased model architecture. In the main paper, we used a biased model with the same architecture as the teacher models (i.e., ResNet-50). We experiment with different network backbones for the biased models in AGRE-KD. The results below do not show significant differences across biased model choices, demonstrating the robustness of the proposed method to the choice of biased model. These results suggest that the gradient direction of any biased pretrained model can provide sufficient guidance for debiased distillation.
@@
Models Architecture KD Debiased Teacher ERM Teacher
Avg WGA Avg WGA
Teacher ViT $93.9_{\pm 0.48}$ $85.0_{\pm 1.46}$ $95.8_{\pm 0.00}$ $38.8_{\pm 0.78}$
ResNet-50 $93.2_{\pm 0.92}$ $90.1_{\pm 0.86}$ $95.8_{\pm 0.17}$ $42.2_{\pm 3.14}$
ConvNeXtV2 $92.9_{\pm 0.07}$ $86.9_{\pm 0.39}$ $95.9_{\pm 0.05}$ $44.4_{\pm 0.78}$
RANDOM ResNet-18 $94.1_{\pm 0.48}$ $79.4_{\pm 3.14}$ $95.9_{\pm 0.07}$ $36.3_{\pm 0.39}$
AVER $94.1_{\pm 0.38}$ $80.0_{\pm 3.10}$ $95.9_{\pm 0.09}$ $36.1_{\pm 0.00}$
AE-KD $93.9_{\pm 0.19}$ $82.2_{\pm 1.57}$ $95.9_{\pm 0.10}$ $40.2_{\pm 1.17}$
AGRE-KD $90.4_{\pm 0.69}$ $87.6_{\pm 0.94}$ $95.9_{\pm 0.03}$ $41.2_{\pm 1.57}$
Table 8: Results with heterogeneous teacher ensemble on the CelebA dataset. We train teachers with different architectures (ViT (base), ResNet-50, and ConvNeXtV2 (base)) and perform knowledge distillation to a ResNet-18 student model.